OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

The slope of the given line is \[ m=\frac 2 3 \]

OpenStudy (anonymous):

The line perpendicular to is has an equation\[ y= -\frac 32 x +b \]

OpenStudy (anonymous):

Find b so the line passes through (-8,2)

OpenStudy (anonymous):

\[ 2 = -\frac 3 2 (-8) + b\\ 2= 12 + b\\ b=-10 \]

OpenStudy (anonymous):

The equation is \[ y= -\frac 3 2 x -10 \]